If karen were two years older than she is she would be twice as old as larry who is eight years younger than she. how old is each one?
Answer by ankor@dixie-net.com(22740)

If karen were two years older than she is, she would be twice as old as larry,
who is eight years younger than she. how old is each one?
:
K = Karen's age now
L = Larry's age now
:
K + 2 = 2L
but it says: L = (K-8), substitute (k-8) for L
K + 2 = 2(K - 8)
:
K + 2 = 2k - 16
:
k - 2k = -16 - 2
:
-k = -18
:
K = 18 is Karen's age now
then
18 - 8 = 10 is Larry's age now
;
;
Check solution in 1st equation
18 + 2 = 2(10)